# Cross-Document Event-Keyed Summarization

### v1

@@ Results
are shown in the right half of Table 2 and are qualitatively similar to those for the report task, with the fine-tuned models generally showing the best overall numbers ($\textbf{R}_{1,2,L}$, CR) or nearly so (BS), although GPT-4o obtains the highest scores on BS and Claude Sonnet on A and F. Once again, nearly all models outperform the report baseline across the board (ZS GPT-4o Mini excepted), . Finally, we note that results on most metrics are much lower in absolute terms compared to the corresponding results from §4.2, testifying to the greater difficulty of the cross-document task.
@@ 4.4 Impact of Extraction Quality
The preceding experiments conditioned summary generation on gold event annotations in order to focus model comparisons narrowly on summarization ability. In most practical settings, however, a (CD)EKS model would instead have to rely on the imperfect outputs of an event extraction model. To probe robustness to extraction errors in a controlled manner, we apply variable amounts of noise to the gold event annotations and evaluate model performance on the resulting examples. Concretely, for each role $R$ of each event, we edit $R$’s arguments with probability $p$. If a role is selected for editing, we then make one of the following edits with equal probability:
@@
    Insert: A new span from the text is added to the argument list for $R$.
@@
    Delete: An argument span is removed at random from the argument list.
@@
    Replace: An argument span is replaced at random with a different span from the text.
For the cross-document task, we apply these edits to the event annotations for both the report and the source. We sample the edits to be made and then prompt an LLM (GPT-4o) to apply them by supplying in a prompt: (1) the text (report or source), (2) the (JSON-formatted) report or source event annotations, and (3) instructions for the edits to be made, generated automatically by populating templatic statements based on the edits sampled. The LLM is free to select an appropriate new span to be used for the Insert and Replace operations. We consider $p\in\{0.1,0.2,0.3,0.4,0.5\}$, using the same sampled edits for all models for a given $p$.
@@ Results
for both tasks are in Table 3. For all models, we observe (near-)monotonic drops in performance for most metrics as $p$ increases. While performance drops are sizable in some cases, they are arguably less radical than one might expect, given the destructiveness of the changes at $p=0.5$, where roughly half of all roles contain extraction errors. This is especially evident in the results for Claude Haiku and GPT-4o Mini on the cross-document task, where (e.g.) $\textbf{R}_{1,2,L}$ scores decrease by only about 3 points from $p=0$ to $p=0.5$, BS by less than 1, and F showing no drop at all. Further, losses on CR (the most explicit measure of extraction ability) are only $\sim$5 points for GPT-4o Mini and $\sim$3 points for Claude Haiku.
@@
4.5 Input Ablations
Following Gantt et al. (2024), Appendix F considers ablations on the input for both tasks, in which we omit the annotated events (Text Only) or the texts (Event Only), and condition summary generation on the ablated inputs. We also present a novel third ablation that omits the annotated arguments only, leaving intact information about the target frame and roles (Text+Schema). Consistent with Gantt et al.’s results, we find that both the text and the full event annotations are needed to obtain the best results (see Tables 6 and 7), indicating that the SEAMuS tasks are not reducible to standard summarization (Text Only), structure-to-text (Event Only), or even a (simpler) hybrid objective (Text+Schema).
@@ 5 Human Evaluation
Lastly, we conduct a human evaluation of the reference and model-generated summaries. We focus our evaluation on the combined task only, comparing the summaries generated by models presented in Table 2 (RB excluded). For the GPT and Claude models, we use the summaries generated in the few-shot setting only (FS), owing to their superior results above. We randomly sampled 30 test set examples and presented the model-generated summaries for these examples, along with the references, to 3 human raters—all English-speaking NLP researchers who did not participate in other parts of this work. Each rater provided a single quality score for each summary based on the ordered list of attributes used by Gantt et al. (2024): factuality, adequacy, coherence, relevancy, and fluency. Scores were given from 1 (low) to 5 (high), with half points allowed. Each rater thus provided $30\times(7+1)\times 1=240$ judgments. Raters were not shown which model produced which summary, and summary presentation order was randomized. [^19] See Appendix D for further details.
@@ Results
Looking at intra-rater distributions, however, it’s unclear how robust these preferences are. Using Wilcoxon rank-sum tests to evaluate pairwise differences in each rater’s scores for a given pair of models, we find that some of these preferences are reliable at $\alpha=0.05$ (e.g. GPT-4o $>$ T5 with $p=0.016$ for the first rater), but none holds up when applying the Bonferroni correction to account for multiple comparisons. We take these results to indicate that our baselines are broadly effective at producing clear and accurate summaries, and that while they may somewhat differentiate themselves on individual metrics, the best models on a more holistic picture may come down to user preference, and there may not be definitive bests even at this scope. This plurality of solid modeling options is encouraging, and suggests flexibility in the application of CDEKS to a range of use cases.
@@ 6 Conclusion
This work has extended the task of event-keyed summarization (EKS) to the cross-document setting (CDEKS). To enable this, we provided an expert reannotation of the FAMuS CDAE dataset, yielding high-quality event argument annotations on all 1,253 examples. We then leveraged these improved annotations to construct SEAMuS—a collection of single- (report) and cross-document summaries on top of FAMuS, further annotating the summaries themselves for event arguments (§3). We benchmarked SEAMuS on a diverse set of baselines, including smaller fine-tuned models, as well as zero- and few-shot prompted LLMs (§4.2, §4.3). We then presented more detailed analysis, assessing the impact of degraded event extraction on summary quality (§4.4), and conducting a set of ablations on the input (§4.5)—concluding with a human evaluation of summary quality (§5). We release SEAMuS, along with our baseline results, to facilitate further work on EKS in both the single- and cross-document settings.
@@
One limitation of this work is the size of the resource. SEAMuS contains 1,253 examples, which, although sufficient for fine-tuning smaller models and for conducting prompting experiments with larger ones, is likely not sufficient for substantive fine-tuning of much larger models.
A second limitation is that the cross-document setting considers only two documents per example. This constraint was imposed by the choice of the FAMuS dataset as the basis for SEAMuS, as cross-document argument annotations in the former were provided only for pairs of report and source texts. Future work expanding the set of source texts would be valuable, and would allow both for richer summaries and for more robust evaluation of models’ ability to accurately synthesize information across possibly differing accounts of events.
Addressing either limitation may require relaxing data quality goals—relying on either crowdsourcing or LLM-powered annotation techniques—as scaling our annotation procedure to substantially more examples or source documents would demand considerable resources.
@@ Appendix A Additional Examples
Below, we show a few examples of the report summaries and their corresponding cross-document summaries to illustrate how the latter are able to provide greater detail about an event of interest relative to the former. We note, however, that this is not always the case: sometimes the source document offers no additional information about the event beyond what is contained in the report, and so the resulting cross-document summary is not able to provide any more detail than the report summary.
@@ Appendix B Training and Evaluation
We use the implementations of ROUGE ($\textbf{R}_{1,2,L}$) and BERTScore (BS) provided by the HuggingFace evaluate library. We implement CEAF-REE (CR) and its soft-match variant (see Tables 6, 7) using the metametric package (Chen et al., 2023). We use the implementation of AlignScore released by the metric’s authors (Zha et al., 2023). [^20] https://github.com/yuh-zha/AlignScore. Lastly, for FActScore, we use the few-shot examples from Wanner et al. (2024) for decomposition and use Llama3.1-8B Instruct (Touvron et al., 2023; Dubey et al., 2024) for both atomic fact decomposition and verification.
@@
The input format for the combined task is:
@@ E.1 License
We release SEAMuS and our code under a CC-BY-SA-4.0 license. As noted in the Ethics section, we intend SEAMuS for academic/research use only.
E.2 Additional Statistics

### v2

@@ Results
are shown in the right half of Table 2 and are qualitatively similar to those for the report task, with the fine-tuned models generally showing the best overall numbers ($\textbf{R}_{1,2,L}$, CR) or nearly so (BS), although GPT-4o obtains the highest scores on BS and Claude Sonnet on A and F. Once again, nearly all models outperform the report baseline across the board (ZS and GPT-4o Mini excepted). Finally, we note that results on most metrics are much lower in absolute terms compared to the corresponding results from §4.2, testifying to the greater difficulty of the cross-document task.
4.4 Input Ablations
Following Gantt et al. (2024), Appendix F considers ablations on the input for both tasks, in which we omit the annotated events (Text Only) or the texts (Event Only), and condition summary generation on the resulting ablated inputs. We also present a novel third ablation that omits the annotated arguments, but leaves intact information about the target frame and roles (Text+Schema). Consistent with Gantt et al.’s results, we find that both the text and the full event annotations are needed to obtain the best results (see Tables 8 and 9 in Appendix F), indicating that the SEAMuS tasks are not reducible to standard summarization (Text Only), structure-to-text (Event Only), or even a (simpler) hybrid objective (Text+Schema).
@@ 4.5 Impact of Extraction Quality
Given that models require event arguments in the input to produce the best summaries (§4.4), a natural next question concerns the sensitivity of these models to noise in the arguments. In real-world scenarios, one generally will not have access to gold arguments (as in §4.2-4.3) and must instead rely on the outputs of an event extraction model.
To probe robustness to extraction errors in a controlled manner, we apply variable amounts of noise to the gold event annotations and evaluate model performance on the resulting inputs. Concretely, for each role $R$ of each event, we edit $R$’s arguments with probability $p$. If a role is selected for editing, we then make one of the following edits with equal probability:
@@
    Insert: A new (incorrect) span from the text is added to the argument list for $R$.
@@
    Delete: An argument span is removed at random from the argument list for $R$.
@@
    Replace: An argument span is replaced at random with an incorrect span from the text.
For the cross-document task, we apply these edits to the event annotations for both the report and the source. We sample the edits to be made uniformly and then prompt an LLM (GPT-4o) to apply them by supplying in a prompt: (1) the text (report or source), (2) the (JSON-formatted) report or source event annotations, and (3) instructions for the edits to be made, generated automatically by populating templatic statements based on the edits sampled. The LLM is free to select an appropriate new span to be used for the Insert and Replace operations. We consider $p\in\{0.1,0.2,0.3,0.4,0.5\}$, using the same sampled edits for all models for a given $p$.
@@ Results
for both tasks are in Table 3. For all models, we observe (near-)monotonic drops in performance for most metrics as $p$ increases. While performance drops are sizable in some cases, they are arguably less radical than we might expect, given the destructiveness of the changes at $p=0.5$, where roughly half of all roles contain extraction errors. This is especially evident in the results for Claude Haiku and GPT-4o Mini on the cross-document task, where (e.g.) $\textbf{R}_{1,2,L}$ scores decrease by only about 3 points from $p=0$ to $p=0.5$, BS by less than 1, and F showing no drop at all. Further, losses on CR (the most explicit measure of extraction ability) are only $\sim$5 points for GPT-4o Mini and $\sim$3 points for Claude Haiku.
@@ 5 Human Evaluation
Lastly, we conduct a human evaluation of the reference and model-generated summaries. We focus our evaluation on the cross-document task, comparing the summaries generated by models presented in Table 2 (excluding RB). For the GPT and Claude models, we use the FS (few-shot) summaries only, owing to their superiority over the ZS results. We randomly sampled 30 test set examples and presented the 7 model-generated summaries for these examples, along with the references, to 3 human raters—all English-speaking NLP researchers who did not participate in other parts of this work. Each rater provided a single quality score for each summary based on the ordered list of attributes used by Gantt et al. (2024): factuality, adequacy, coherence, relevancy, and fluency. Scores were given from 1 (low) to 5 (high), with half points allowed. Each rater thus provided $30\times(7+1)\times 1=240$ judgments. Raters were not shown which model produced which summary, and summary presentation order was randomized. [^18] See Appendix D for further details.
@@ Results
[…]
@@ 6 Conclusion
This work has extended the task of event-keyed summarization (EKS) to the cross-document setting (CDEKS). To enable this, we provided an expert reannotation of the FAMuS CDAE dataset, yielding high-quality event argument annotations on all 1,265 examples. We then leveraged these improved annotations to construct SEAMuS—a collection of single- (report) and cross-document summaries on top of FAMuS, further annotating the summaries themselves for event arguments (§3). We benchmarked SEAMuS on a diverse set of baselines, including smaller fine-tuned models, as well as zero- and few-shot prompted LLMs (§4.2, §4.3). We then presented more detailed analysis, conducting a comprehensive set of input abalations (§4.4), assessing the impact of degraded event extraction on summary quality (§4.5), and finally concluding with a human evaluation of summary quality (§5). We release SEAMuS, along with our baseline results, to facilitate further work on EKS in both the single- and cross-document settings.
@@
One limitation of this work is SEAMuS’s size: 1,265 examples is sufficient for fine-tuning smaller models and for conducting prompting experiments with larger ones, but is likely insufficient for substantive fine-tuning of very large models.
A second limitation is that the cross-document setting considers only two documents per example. This constraint was imposed by the choice of the FAMuS dataset as the basis for SEAMuS, as cross-document argument annotations in the former were provided only for pairs of report and source texts. Future work expanding the set of source texts would be valuable, and would allow both for richer summaries and for more robust evaluation of models’ ability to accurately synthesize information across possibly differing accounts of events (cf. Huang et al. (2024)), as information conflicts are more common outside of Wikipedia citations.
We note, however, that addressing either limitation may require relaxing data quality standards—relying on crowdsourcing or LLM-powered annotation techniques—as scaling our annotation procedure to many more examples or source documents would demand considerable resources. It was only thanks to the above restrictions that we were able to provide expert annotations for SEAMuS.
@@ Appendix A Additional Examples
Below, we show a few examples of the report summaries and their corresponding cross-document summaries to illustrate how the latter typically provide greater detail about an event of interest relative to the former. We note, however, that this is not always the case: sometimes the source document offers no additional information about the event beyond what is contained in the report.
@@ Appendix B Training and Evaluation
We use the implementations of ROUGE ($\textbf{R}_{1,2,L}$) and BERTScore (BS) provided by the HuggingFace evaluate library. We implement CEAF-REE (CR) and its soft-match variant (see Tables 8, 9) using the metametric package (Chen et al., 2023). We use the implementation of AlignScore released by the metric’s authors (Zha et al., 2023). [^20] https://github.com/yuh-zha/AlignScore. Lastly, for FActScore, we use the few-shot examples from Wanner et al. (2024) for decomposition and use Llama3.1-8B Instruct (Touvron et al., 2023; Dubey et al., 2024) for both atomic fact decomposition and verification.
@@
The input format for the cross-document task is:
@@ E.1 License
We release SEAMuS and our code under a CC-BY-SA-4.0 license. As noted in the Ethics section, we intend SEAMuS for research use only, not for commercial purposes.
E.2 Additional Summary Statistics
@@
E.3 Inter-Annotator Agreement
As we note in §3, there was no redundancy in the SEAMuS annotation process: corrections to the FAMuS arguments, writing of summaries, and annotation of summary arguments were performed by a single annotator for each example. However, annotators did conduct a 10-example practice annotation for both the report and cross-document tasks. Thus, to give some (limited) sense of the inter-annotator agreement, Table 5 and Table 6 present pairwise comparisons of annotators’ annotations on these 10 items for the report and cross-document tasks (respectively) using the reference-based metrics from Table 2 (plus the edit distance version of CR, CRsoft; see Appendix F). We treat annotations produced by annotators in the $P$ column as “predictions” to be evaluated against the “reference” annotations produced by annotators in the $R$ column. Two important notes:
  1. 1.
    Because all of these metrics are $\text{F}_{1}$ scores, the distinction between $P$ and $R$ is moot and reversing $P$ and $R$ for any given pair would yield the same results. In both tables, we report results for all unordered annotator pairs, as well as the average across all pairs.
  2. 2.
    Because these were practice annotations, none of them were included in the final SEAMuS dataset. We would thus expect the numbers reported here to be an underestimate of the level of agreement on the main task, had we had redundancy.
$P$ $R$ $\textbf{R}_{1}$ $\textbf{R}_{2}$ $\textbf{R}_{L}$ BS CR $\textbf{CR}_{\text{soft}}$ A1 A2 $67.1$ $44.5$ $52.6$ $93.4$ $72.7$ $86.9$ A1 A3 $72.4$ $53.9$ $63.1$ $94.4$ $76.4$ $86.5$ A2 A3 $77.2$ $60.5$ $62.5$ $94.7$ $75.9$ $89.6$ Avg. $72.2$ $53.0$ $59.2$ $94.2$ $75.0$ $87.7$
Table 5: Inter-annotator agreement on the 10 practice examples from the SEAMuS report summary annotation, as given by the reference-based metrics we report in §4, treating annotator $P$’s responses as predictions and $R$’s responses as references (the reverse is equivalent, since these metrics are symmetric).
$P$ $R$ $\textbf{R}_{1}$ $\textbf{R}_{2}$ $\textbf{R}_{L}$ BS CR $\textbf{CR}_{\text{soft}}$ A1 A2 $64.8$ $42.4$ $53.3$ $94.0$ $40.5$ $56.6$ A1 A3 $64.7$ $43.8$ $51.6$ $93.0$ $50.6$ $65.0$ A1 A4 $45.8$ $22.7$ $33.0$ $90.7$ $49.8$ $65.2$ A1 A5 $69.0$ $48.1$ $58.8$ $94.6$ $46.8$ $62.8$ A2 A3 $77.7$ $66.8$ $72.4$ $94.9$ $50.6$ $65.6$ A2 A4 $55.4$ $33.9$ $40.8$ $91.2$ $50.7$ $66.8$ A2 A5 $72.0$ $56.1$ $64.9$ $94.1$ $50.6$ $66.9$ A3 A4 $55.4$ $33.2$ $41.9$ $90.7$ $51.8$ $69.0$ A3 A5 $71.0$ $57.2$ $61.8$ $93.6$ $52.6$ $69.8$ A4 A5 $48.2$ $27.2$ $37.3$ $90.7$ $52.8$ $69.6$ Avg. $62.4$ $43.1$ $51.6$ $92.8$ $49.7$ $65.7$
Table 6: Inter-annotator agreement on the 10 practice examples from the SEAMuS cross-document summary annotation, as given by the reference-based metrics we report in §4, treating annotator $P$’s responses as predictions and $R$’s responses as references (the reverse is equivalent, since these metrics are symmetric).
@@ F.2 Input Ablations
F.3 Argument Recovery by Role
Table 12 and Table 13 show CR and CRsoft results (respectively) on the cross-document task broken down by role for the 20 roles with highest support (number of annotated arguments) in the SEAMuS training split.
Comparing the tables reveals an interesting dichotomy. For CR, no model is consistently dominant across all roles, with fine-tuned models collectively obtaining the best results on 12 of the 20 and few-shot prompted models obtaining the best results on the remaining 8. The CRsoft results, by contrast, heavily favor GPT-4o, which achieves the best scores on 13 roles. Here, the fine-tuned models are top-performing on only 4 roles.
We believe the same factor discussed in subsection F.2 explains this dichotomy: whereas CR requires exact span match—and thus will tend to favor models able to learn span boundary conventions through fine-tuning—CRsoft does not, and rewards spans proportional to their edit distance from the reference. Thus, CRsoft reveals the LLMs (and GPT-4o above all) to be effective in producing summaries that recover the correct arguments, albeit with more lexical modifications relative to the reference.
[…]
Table 12: CR $\text{F}_{1}$ results on test set cross-document summaries for the top 20 roles with highest support (# arguments) in the SEAMuS training split (which has 3,004 total arguments). Results with GPT and Claude models are from the few-shot (FS) setting. Best results for each role are bolded.
Role Support GPT-4o M GPT-4o Claude H Claude S BART PEGASUS T5 Time $523$ $57.37$ $65.22$ $49.22$ $50.52$ $60.54$ $61.40$ $\mathbf{65.43}$ Place $499$ $44.84$ $\mathbf{53.39}$ $37.46$ $37.67$ $46.93$ $47.67$ $52.15$ Agent $240$ $65.40$ $\mathbf{67.47}$ $59.78$ $50.16$ $66.72$ $57.99$ $62.06$ Theme $94$ $73.21$ $\mathbf{73.96}$ $69.22$ $72.96$ $71.75$ $64.71$ $68.59$ Entity $65$ $66.33$ $\mathbf{76.13}$ $63.16$ $60.07$ $63.34$ $63.12$ $69.44$ Patient $53$ $\mathbf{76.28}$ $73.52$ $74.41$ $68.62$ $70.88$ $73.28$ $73.36$ Goal $49$ $50.53$ $\mathbf{66.30}$ $50.19$ $45.42$ $47.30$ $53.63$ $60.30$ Event $43$ $52.75$ $\mathbf{63.52}$ $52.76$ $42.95$ $45.97$ $52.09$ $35.38$ Cause $42$ $47.66$ $\mathbf{52.99}$ $39.41$ $42.03$ $43.23$ $43.87$ $49.82$ Experiencer $39$ $68.83$ $\mathbf{91.48}$ $82.86$ $69.44$ $56.03$ $60.99$ $61.06$ Victim $39$ $62.13$ $\mathbf{74.79}$ $71.75$ $60.13$ $65.62$ $60.53$ $55.89$ Goods $38$ $42.77$ $\mathbf{79.33}$ $33.35$ $50.68$ $61.94$ $57.46$ $24.39$ Protagonist $38$ $60.13$ $\mathbf{72.52}$ $54.02$ $66.03$ $66.57$ $67.63$ $61.24$ Source $30$ $78.00$ $\mathbf{84.13}$ $60.54$ $65.55$ $66.80$ $55.03$ $69.62$ Topic $26$ $19.06$ $19.06$ $17.75$ $20.42$ $21.70$ $\mathbf{35.80}$ $12.96$ Speaker $25$ $58.21$ $61.41$ $71.85$ $64.91$ $64.92$ $\mathbf{73.03}$ $66.28$ Addressee $22$ $44.60$ $\mathbf{89.49}$ $40.21$ $52.63$ $63.63$ $73.95$ $41.94$ Stimulus $21$ $38.61$ $70.46$ $54.55$ $66.70$ $53.39$ $\mathbf{74.75}$ $57.65$
Table 13: $\textbf{CR}_{\text{soft}}$ (distinct from CR; see §F.2) $\text{F}_{1}$ results on test set cross-document summaries for the top 20 roles with highest support (# arguments) in the SEAMuS training split (which has 3,004 total arguments). Results with the GPT and Claude models are from the few-shot (FS) setting. Best results for each role are bolded.